A population has a mean μ 84 and a standard deviation -27 Find the mean and standard deviation of a sampling dstribution of sample means with sample sze n-81 (Simplity your answer) % [] (Simplify your answer )
Use the given confidence interval to find the margin of error and the sample mean. (13.3,23.1) The sample mean is(Type an integer or a decimal.) in of error is . (Type an integer or a decimal.)
Answer #1

1)

\mux =84

\sigmaxbar =27/sqrt(81)=3

2)

sample mean =(13.3+23.1)/2 =18.2

margin of error =(23.1-13.3)/2=4.9